# Parallel Token Swapping for Qubit Routing

### v1

@@ 1 Introduction
Motivated by recent applications in qubit routing on quantum computers, we study the parallel token swapping problem first introduced by Alon et al. [2]. Here, one is permitted to swap multiple pairs of pairwise disjoint adjacent tokens at the same time. The parallel token swapping problem has been shown to NP-Hard even on subdivided stars [1]; we provide the first known constant factor approximations for several common graph topologies arising as substructures in modern quantum computers [11, 19, 28, 7]. In addition, we compare our approximation factors to the so called stretch factor of a natural lower bound (the maximum geodesic distance of a token between its positions in the initial and final configurations) and show that our approximation factors are tight with respect to this lower bound in some cases. The stretch factor refers to the maximum ratio of $OPT/LB$ over all instances where $OPT$ is an optimal solution and $LB$ is the lower bound.
@@ 1.1 Motivation
Quantum algorithms are usually designed under the assumption that any two qubits in a quantum computer can interact via a quantum gate. However, in practical hardwares, physical qubits are laid out in a two-dimensional or three-dimensional topology and only adjacent qubits can interact via gates. Fortunately, this discrepancy can be handled by using SWAP gates, which switch the positions of two adjacent qubits (we refer the reader to [20] for a detailed discussion). This process is referred to as qubit routing, and is a necessary step for any quantum compiler. SWAP gates should be added carefully, as they increase the size and depth of the quantum algorithm, which adversely affects the fidelity of a quantum circuit (see, for example [10, 11]). Hence, the goal of the qubit routing problem is to find an initial positioning of the qubits and a sequence of SWAP gates that minimizes the size and/or depth of the quantum circuit. The initial allocation subproblem is equivalent to the graph isomorphism problem [25, 6] while the problem of swapping qubits is equivalent to the the token swapping problem [30]. The parallel token swapping problem optimizes the depth of the quantum circuit.
@@
The qubit routing problem has proven to be extremely hard, owing in part to the complexity of the subgraph isomorphism and token swapping problems. Exact algorithms can be found in the works of [25, 19, 29]. These algorithms, however, have highly exponential running times that do not scale well with the size of the quantum computer. Consequently, there is a large body of research on developing empirically good heuristics for the problem [7, 26, 33, 15, 23, 13, 18, 22, 24, 17, 28]. We refer the reader to the following survey on algorithms for the qubit routing problem [3]. Very recently, Wagner at al. [28] considered the problem of optimizing the size of the quantum algorithm and observed that approximation algorithms for the token swapping problem can be leveraged to provide superior heuristics for the qubit routing problem. In particular, the previously known 4-approximation algorithm for the token swapping problem [16] can be used to guide the heuristic when deciding on qubit allocations. They attempted to extend their results to the problem of optimizing the depth of the quantum circuit, but were limited by the lack of good approximation algorithms for the parallel token swapping problem. Their methodology is to use the 4-approximation algorithm for the non-parallel token swapping problem and combine gates into layers sequentially. However, optimal solutions to the parallel token swapping problem can be very different from optimal solutions to the token swapping problem, bringing to light the importance of designing good approximation algorithms for the parallel token swapping problem. In this work, we provide approximation algorithms for the parallel token swapping problem on common graphs arising as substructures in current quantum topologies ([11, 19, 28, 7]). For example IBM’s heavy hex lattice quantum architecture can be viewed as a union of cycle graphs and also as a union of subdivided star graphs.
@@
A complete solution to the qubit routing problem requires a sequence of qubit allocations under which the quantum gates can be implemented, and also a sequence of swaps to move from one allocation to another. The stretch factor of lower bounds for the token swapping problem is a crucial tool linking the two parts. If the stretch factor is reasonably bounded, one can use this lower bound as a proxy to determine good qubit allocations. For instance, Wagner et al. [28] designed a heuristic for the qubit routing problem using the half sum lower bound of Miltzow et al. [16], which was shown to have a stretch factor of 4 for the token swapping problem. They were able to use this lower bound as an objective in the integer program of [19] to develop superior heuristics for the qubit routing problem. Hence, a careful analysis of the stretcth factor of natural linear lower bounds is essential for a better understanding of the qubit routing problem. In the parallel token swapping problem, the natural linear lower bound is the maximum geodesic distance of a token between its initial and final configurations. We study the stretch factor of this lower bound on the graph topologies mentioned earlier.
@@ 1.2 Our Contributions
We prove this by providing two instances of the problem with the same distance vector $(d_{1},\ldots,d_{n})$. The first instance is on the complete graph and it is well known that the optimal solution on the complete graph uses at most two steps. The second instance is on the cycle graph and we argue that any solution must use at least $n-1$ steps. The above result provides evidence that the design of good approximation algorithms for (PTS) must rely on utilizing the structure of the underlying graph. Due to this, we focus our attention to specific graph topologies arising in quantum computers, beginning with a result on cycle graphs. Throughout, we use $OPT$ to denote the optimal solution value of a given instance of (PTS).
@@
    If $h=2$, we can find a solution with value at most $12OPT+$ 1 in polynomial time.
@@
Subdivided Stars Graphs $\mathbf{4OPT+\min\{OPT,h}+1\}$ $\mathbf{\Theta(h)}$ $\mathbf{4OPT+\min\{OPT,h\}+1}$
Ladder Graphs $\mathbf{2OPT+1}$ $\mathbf{2OPT+1}$
@@
The algorithm when $h=2$ works by isolating a spanning path from the grid graph and utilizing the algorithm in [12] to reconfigure the tokens along this line subgraph. The optimal solution however, can utilize the edges in the grid graph, not on this path. We show that one simulate those swaps on the path without spending too many additional steps. This algorithm does not scale well with $h$. Instead, for larger values of $h$, we perform a three phase algorithm where we reconfigure token along rows, then column, and then rows again to obtain the final configuration. At the end of the first phase, we ensure that each column has exactly one token whose position in the final configuration is in row $i$, for any $i$. We then reconfigure the columns to bring each token to its correct row, and finally reconfigure the rows in parallel. The phases along rows can be bounded by the size of the rows ($h$), while the phase along columns can be bounded by our next result on the stretch factor of line graphs.
@@
Next, we extend our results to the colored versions of token swaping problem. The first one is the colored (PTS) , where each token and each vertex is assigned a color, and every token should be routed to a vertex that shares its color. In this setting the number of colors can be strictly less than the number of vertices. A (PTS) problem on a graph with $n$ vertices can be viewed as a colored (PTS) on the same graph with $n$ colors where each token has a unique color.
Finally, we consider a variant of (PTS) when the number of tokens is smaller than the number of nodes in the graph. This happens, for instance, if some qubits are not operational on the quantum computer. In this case, all tokens can be assigned a unique color and all all empty positions can be assigned a common color and dummy tokens with that color can be added in. We refer to this special case of the colored (PTS) as the incomplete (PTS).
@@ 2 Preliminaries
$MC(v)=\begin{cases}C(w),&\text{if }(v,w)\in M\\ C(v),&\text{otherwise}\end{cases}$
@@ 3 The stretch factor
Complete Graph Instance Let $G$ be the complete graph with vertex set $V=\{1,2,\ldots,2r\}=T$. Let $C_{0}$ be the initial configuration where $C_{0}(i)=i+1$ if $i\leq 2r-1$ and $C_{0}(2r)=1$. Consider the following two matchings, $M_{1}=\{(i,2r-i):i=1,\ldots,r-1\}$ and $M_{2}=\{(i,2r-1-i):i=1,\ldots,r-1\}\cup\{(2r-1,2r)\}$. Then, $M_{2}M_{1}C_{0}=Id$. Here, $OPT=2$ and $d_{i}=1$ for any token $i$.
Cycle Graph Instance Let $G$ be a cycle graph with vertex set $V=\{1,2,\ldots,2r\}=T$, and edge set $E=\{(i,i+1):i\leq 2r-1\}\cup\{(2r,1)\}$. Let $C_{0}$ be the initial configuration where $C_{0}(i)=i+1$ if $i\leq 2r-1$ and $C_{0}(2r)=1$. So the goal is to move every token one step counter-clockwise. We argue that any solution must use at least $12r-$ 1 matchings by using the following charging scheme. Initially every token has a charge of zero. If a token moves clockwise, we decrease its charge by one and if a token moves counter-clockwise, we increase its charge by one. In the final configuration, each token has moved one step counter-clockwise and so has a charge of 1 (modulo $2r$). Any matching does not change the total charge summed over all tokens as the number of tokens that move clockwise is equal to the number of tokens that move counter-clockwise. Since the total charge on all the tokens in the final configuration is zero, and every token has a charge of 1 (modulo $2r$), at least one of the tokens must have a negative charge. This charge must be $-2r+1$ or smaller. Additionally, any matching changes the charge on a token by at most ±$\pm 1$ and so we need at least $12r-$ 1 matchings. We also observe that the optimal solution uses exactly $12r-$ 1 matchings. Set $M_{i}=(i,i+1)$ for $i=1,\ldots,2r-1$ and then the sequence of matchings $M_{1},\ldots,M_{2r-1}$ routes $C_{0}$ to $Id$. Here $OPT=n-1$ and $d_{i}=1$ for any token $i$. ∎
@@
Consider the cycle graph instance from the proof above. Here $OPT=n-1$ while $d_{max}=1$. Hence, the stretch factor of the lower bound $d_{max}$ is at least $n-1$. Furthermore, we can ignore an edge of the graph, say $(2r,1)$ and treat the problem as an instance on the line graph. Then, by Observation 9, any instance on the cycle graph can be solved in at most $n$ steps. Hence, the stretch factor of the lower bound $d_{max}$ is at most $n$. ∎
@@ 3.1 Stretch Factor for Line Graphs
In this section, we we analyze the stretch factor for line graphs with respect to the maximum geodesic distance $d_{max}$, and show that this stretch factor is asymptotically tight. Let the underlying graph be a line graph $P_{n}=(V,E)$ where $V=(1,2,\dots,n)$ and $E=\{(i,i+1):1\leq i\leq n-1\}$. With this topology, (PTS) can be viewed as a parallel version of bubble sort. We show that the stretch factor of $d_{max}$ is 2 for line graphs by proving that the odd-even algorithm achieves this bound.
@@
We create a potential function $\Phi_{C}$ for every token in the configuration $C$. Informally, for any token $t$ we want to judge approximately how many iterations it will take for all tokens that must cross $t$ to do so. Although the following description of the potential function is heavy on notation, the underlying ideas are actually not complex. In simple terms, for each possible prefix $[1,j]$ of vertices, we count the number of tokens in the prefix that wish to cross the token at $i$ along with the tokens on vertices in $[j+1,i-1]$ that do not wish to cross the token at $i$, and choose the largest count over all prefixes. After doing the same for all suffixes, combining the two counts yields our potential value.
Now we will define the potential function more formally. Given a vertex $i$ with $C(i)=t$ and some $j<i$, we define the set of prefixes $Q_{C}^{-}(t)$ as the sets of vertices $\{[1,j]:C(i^{\prime})>t\text{ for some }i^{\prime}\in[1,j]\}$ where at least one present token wishes to cross $t$. Similarly, for $j>i$, we define suffixes $Q_{C}^{+}(t)$ as the sets of vertices $\{[j,n]:C(i^{\prime})<t\text{ for some }i^{\prime}\in[j,n]\}$. For a prefix $Q^{-}\in Q_{C}^{-}(t)$ where $Q^{-}=[1,j]$, we let $\overline{Q^{-}}=[j+1,i-1]$, and similarly for a suffix $Q^{+}=[j,n]$, we let $\overline{Q^{+}}=[i+1,j-1]$.
@@
Fix a token $t$ with $C^{-1}(t)=v$. First, suppose that there are tokens on both sides of $t$ that wish to cross it. That is, there are vertices $i<v$ with $C(i)>t$ and $j>v$ with $C(j)<t$. We note that $\Phi_{C}(t)$ counts tokens, and the only tokens we may count lie between (inclusively) the smallest choice of $i$ and the largest choice of $j$. The smallest $i$ can be is $t-d_{max}+1$; any smaller, and there would be a token more than a distance of $d_{max}$ from its destination. Similarly, the largest $j$ can be is $t+d_{max}-1$. There are a total of $12d_{max}-$ 1 tokens in this range, and we may count all but one (the token $t$ itself) as part of $\Phi_{C}(t)$. So $\Phi_{C}(t)\leq 2d_{max}-2$.
@@
If no tokens wish to cross $t$, then the $\Phi_{C}(t)=0$, and the result is trivially true. In all cases, we have that $\Phi_{C}(t)\leq 2d_{max}-1$. ∎

### v2

@@ 1 Introduction
Motivated by recent applications in qubit routing on quantum computers, we study the parallel token swapping problem first introduced by Alon et al. [2]. Here, one is permitted to swap multiple pairs of pairwise disjoint adjacent tokens at the same time. The parallel token swapping problem has been shown to be NP-Hard even on subdivided stars [1]; we provide the first known constant factor approximations for several common graph topologies arising as substructures in modern quantum computers [13, 21, 30, 8]. In addition, we compare our approximation factors to the so called stretch factor of a natural lower bound (the maximum geodesic distance of a token between its positions in the initial and final configurations) and show that our approximation factors are tight with respect to this lower bound in some cases. The stretch factor refers to the maximum ratio of $OPT/LB$ over all instances where $OPT$ is an optimal solution and $LB$ is the lower bound.
@@ 1.1 Motivation
Quantum algorithms are usually designed under the assumption that any two qubits in a quantum computer can interact via a quantum gate. However, in practical hardware, physical qubits are laid out in a two-dimensional or three-dimensional topology and only adjacent qubits can interact via gates. Fortunately, this discrepancy can be handled by using SWAP gates, which switch the positions of two adjacent qubits (we refer the reader to [22] for a detailed discussion). This process is referred to as qubit routing, and is a necessary step for many quantum compilers. SWAP gates should be added carefully, as they increase the size and depth of the quantum algorithm, which adversely affects the fidelity of a quantum circuit (see, for example [12, 13]). Hence, the goal of the qubit routing problem is to find an initial positioning of the qubits and a sequence of SWAP gates that minimizes the size and/or depth of the quantum circuit. The initial allocation subproblem is equivalent to the graph isomorphism problem [27, 7] while the problem of swapping qubits is equivalent to the the token swapping problem [32]. The parallel token swapping problem optimizes the depth of the quantum circuit.
@@
The qubit routing problem has proven to be extremely hard, owing in part to the complexity of the subgraph isomorphism and token swapping problems. Exact algorithms can be found in the works of [27, 21, 31]. These algorithms, however, have highly exponential running times that do not scale well with the size of the quantum computer. Consequently, there is a large body of research on developing empirically good heuristics for the problem [8, 28, 35, 17, 25, 15, 20, 24, 26, 19, 30]. We refer the reader to the following survey on algorithms for the qubit routing problem [4]. Very recently, Wagner at al. [30] considered the problem of optimizing the size of the quantum algorithm and observed that approximation algorithms for the token swapping problem can be leveraged to provide superior heuristics for the qubit routing problem. In particular, the previously known 4-approximation algorithm for the token swapping problem [18] can be used to guide the heuristic when deciding on qubit allocations. They attempted to extend their results to the problem of optimizing the depth of the quantum circuit, but were limited by the lack of good approximation algorithms for the parallel token swapping problem. Their methodology is to use the 4-approximation algorithm for the non-parallel token swapping problem and combine gates into layers sequentially. However, optimal solutions to the parallel token swapping problem can be very different from optimal solutions to the token swapping problem, bringing to light the importance of designing good approximation algorithms for the parallel token swapping problem. In this work, we provide approximation algorithms for the parallel token swapping problem on common graphs arising as substructures in current quantum topologies ([13, 21, 30, 8]). For example IBM’s heavy hex lattice quantum architecture can be viewed as a union of cycle graphs and also as a union of subdivided star graphs. We note that designing good algorithms for more complex graph topologies (including IBM’s heavy hex and Google’s diagonal grid) using algorithms for its subgraphs provided in this paper requires further research.
@@
A complete solution to the qubit routing problem requires a sequence of qubit allocations under which the quantum gates can be implemented, and also a sequence of swaps to move from one allocation to another. The stretch factor of lower bounds for the token swapping problem is a crucial tool linking the two parts. If the stretch factor is reasonably bounded, one can use this lower bound as a proxy to determine good qubit allocations. For instance, Wagner et al. [30] designed a heuristic for the qubit routing problem using the half sum lower bound of Miltzow et al. [18], which was shown to have a stretch factor of 4 for the token swapping problem. They were able to use this lower bound as an objective in the integer program of [21] to develop superior heuristics for the qubit routing problem. Hence, a careful analysis of the stretch factor of natural linear lower bounds is essential for a better understanding of the qubit routing problem. In the parallel token swapping problem, the natural linear lower bound is the maximum geodesic distance of a token between its initial and final configurations. We study the stretch factor of this lower bound on the graph topologies mentioned earlier.
We should note that qubit routing is also possible without SWAP gates. In particular, there are Hamiltonian routing methods that are stronger than SWAP based routing. There are also methods based on local operations, measurement and feedback called LOCC that are akin to qubit teleportation. We refer the reader to [3] and [10] for discussions on the advantages and disadvantages of these methods. These other routing methods are not considered in this paper.
@@ 1.2 Our Contributions
We prove this by providing two instances of the problem with the same distance vector $(d_{1},\ldots,d_{n})$. The first instance is on the complete graph and it is well known that the optimal solution on the complete graph uses at most two steps. The second instance is on the cycle graph and we argue that any solution must use at least $n-1$ steps. The above result suggests that the design of good approximation algorithms for (PTS) must exploit the topology of the underlying graph. Due to this, we focus our attention to specific graph topologies arising in quantum computers, beginning with a result on cycle graphs. Throughout, we use $OPT$ to denote the optimal solution value of a given instance of (PTS).
@@
    If $h=2$, we can find a solution with value at most $22OPT+$ 2 in polynomial time.
@@
Subdivided Stars Graphs $\mathbf{4OPT+\min\{OPT,h\}+1}$ $\mathbf{\Theta(h)}$ $\mathbf{4OPT+\min\{OPT,h\}+1}$
Ladder Graphs $\mathbf{2OPT+2}$ $\mathbf{2OPT+2}$
@@
The algorithm when $h=2$ works by isolating a spanning path from the grid graph and utilizing the algorithm in [14] to reconfigure the tokens along this line subgraph. The optimal solution however, may utilize the edges in the grid graph that are not on this path. We show that those swaps can be simulated by swaps on the path without spending too many additional steps. This algorithm does not scale well with $h$. Instead, for larger values of $h$, we perform a three phase algorithm where we reconfigure token along rows, then column, and then rows again to obtain the final configuration. At the end of the first phase, we ensure that each column has exactly one token whose position in the final configuration is in row $i$, for any $i$. We then reconfigure the columns to bring each token to its correct row, and finally reconfigure the rows in parallel. The phases along rows can be bounded by the size of the rows ($h$), while the phase along columns can be bounded by our next result on the stretch factor of line graphs.
@@
Next, we extend our results to the colored versions of token swapping problem. The first one is the colored (PTS) , where each token and each vertex is assigned a color, and every token should be routed to a vertex that shares its color. In this setting the number of colors can be strictly less than the number of vertices. A (PTS) problem on a graph with $n$ vertices can be viewed as a colored (PTS) on the same graph with $n$ colors where each token has a unique color.
Finally, we consider a variant of (PTS) when the number of tokens is smaller than the number of nodes in the graph. This happens, for instance, if some qubits are not operational on the quantum computer. In this case, all tokens can be assigned a unique color and all empty positions can be assigned a common color and dummy tokens with that color can be added in. We refer to this special case of the colored (PTS) as the incomplete (PTS).
@@ 2 Preliminaries
$MC(v)=\begin{cases}C(w),&\text{if }(v,w)\in M\\ C(v),&\text{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}otherwise.}}}\end{cases}$
@@
Remark 7.
Note that for the purposes of minimizing the size of the sequence of matchings $k$, we can assume that no pair of consecutive matchings $(M_{i},M_{i+1})$ shares a common edge $e=(u,v)$. Indeed if $e\in M_{i}\cap M_{i+1}$, then $M_{i+1}M_{i}C^{\prime}=M^{\prime}_{i+1}M^{\prime}_{i}C^{\prime}$ for any configuration $C^{\prime}$ where $M^{\prime}_{i+1}=M_{i+1}\setminus\{e\}$ and $M^{\prime}_{i}=M_{i}\setminus\{e\}$. Hence, without loss of generality, we shall assume throughout this paper that for any sequence of matchings $M_{i}\cap M_{i+1}=\emptyset$.
@@ 3 The stretch factor
We will define both graphs on the vertex set $V=\{1,2,\dots,2r\}=T$, where the only difference in the two graphs is the edge set. In both cases, we will let $C_{0}$ be the initial configuration where $C_{0}(i)=i-1$ for $2\leq i\leq 2r$ and $C_{0}(1)=2r$.
$v_{3}$$t_{2}$$v_{4}$$t_{3}$$v_{5}$$t_{4}$$v_{6}$$t_{5}$$v_{7}$$t_{6}$$v_{8}$$t_{7}$$v_{1}$$t_{8}$$v_{2}$$t_{1}$diametrically oppositeclockwise
Figure 1: Orientation of Cycle Graphs
Cycle Graph Instance Let $G$ be a cycle graph on $V$ with edge set $E=\{(i,i+1):1\leq i\leq 2r-1\}\cup\{(2r,1)\}$. For ease of explanation, we will say that the vertices in $V$ are arranged in a circular manner, where vertex $i+1$ is clockwise from vertex $i$ (for $11\leq i\leq 2r-$ 1) and vertex $1$ is clockwise from vertex $2r$. Then the goal is to move every token one step counter-clockwise. This configuration is depicted in Figure 1 where $v_{i}$ refers to vertex $i$ and $t_{i}$ refers to token $i$. We use this notation in all figures throughout the paper. We argue that any solution must use at least $12r-$ 1 matchings by using the following charging scheme. Initially every token has a charge of zero. If a token moves clockwise, we decrease its charge by one and if a token moves counter-clockwise, we increase its charge by one. In the final configuration, each token has moved one step counter-clockwise and so has a charge of 1 (modulo $2r$). Any matching does not change the total charge summed over all tokens as the number of tokens that move clockwise is equal to the number of tokens that move counter-clockwise. Since the total charge on all the tokens in the final configuration is zero, and every token has a charge of 1 (modulo $2r$), at least one of the tokens must have a negative charge. This charge must be $-2r+1$ or smaller. Additionally, any matching changes the charge on a token by at most ±$\pm 1$ and so we need at least $12r-$ 1 matchings. We also observe that the optimal solution uses exactly $12r-$ 1 matchings. Set $M_{i}=(i,i+1)$ for $i=1,\ldots,2r-1$ and then the sequence of matchings $M_{1},\ldots,M_{2r-1}$ routes $C_{0}$ to $Id$. Here $OPT=n-1$ and $d_{i}=1$ for any token $i$.
Complete Graph Instance Let $G$ be the complete graph on vertex set $V$. Consider the following two matchings, $M_{1}=\{(i,2r-i):i=1,\ldots,r-1\}$ and $M_{2}=\{(i,2r-1-i):i=1,\ldots,r-1\}\cup\{(2r-1,2r)\}$. Then, $M_{2}M_{1}C_{0}=Id$. Here, $OPT=2$ and $d_{i}=1$ for any token $i$ (see Figure 2). ∎
[…]
$v_{3}$$t_{3}$$v_{4}$$t_{4}$$v_{5}$$t_{5}$$v_{6}$$t_{6}$$v_{7}$$t_{7}$$v_{8}$$t_{8}$$v_{1}$$t_{1}$$v_{2}$$t_{2}$(c) $M_{2}M_{1}C_{0}=Id$
Figure 2: Qubit Routing on a complete graph using two matchings.
@@
Consider the cycle graph instance from the proof above. Here $OPT=n-1$ while $d_{max}=1$. Hence, the stretch factor of the lower bound $d_{max}$ is at least $n-1$. Furthermore, we can ignore an edge of the graph, say $(2r,1)$ and treat the problem as an instance on the line graph. Then, by Observation 10, any instance on the cycle graph can be solved in at most $n$ steps. Hence, the stretch factor of the lower bound $d_{max}$ is at most $n$. ∎
@@ 3.1 Stretch Factor for Line Graphs
In this section, we analyze the stretch factor for line graphs with respect to the maximum geodesic distance $d_{max}$, and show that this stretch factor is asymptotically tight. Let the underlying graph be a line graph $P_{n}=(V,E)$ where $V=\{1,2,\dots,n\}$ and $E=\{(i,i+1):1\leq i\leq n-1\}$. With this topology, (PTS) can be viewed as a parallel version of bubble sort. We show that the stretch factor of $d_{max}$ is 2 for line graphs by proving that the odd-even algorithm achieves this bound.
@@
We create a potential function $\Phi_{C}$ for every token in the configuration $C$. Informally, for any token $t$ we want to judge approximately how many iterations are required to arrive at a configuration in which no tokens wish to cross $t$. Although the following description of the potential function is heavy on notation, the underlying ideas are actually not complex. In simple terms, for each possible prefix $[1,j]$ of vertices, we count the number of tokens in the prefix that wish to cross the token at $i$ along with the tokens on vertices in $[j+1,i-1]$ that do not wish to cross the token at $i$, and choose the largest count over all prefixes. After doing the same for all suffixes, combining the two counts yields our potential value.
Now we will define the potential function more formally. Given a vertex $i$ with $C(i)=t$, we define the set of prefixes $Q_{C}^{-}(t)$ as the sets of vertices $\{[1,j]:C(i^{\prime})>t\text{ for some }i^{\prime}\in[1,j],j<i\}$ where at least one present token wishes to cross $t$. Similarly, we define the set of suffixes $Q_{C}^{+}(t)$ as the sets of vertices $\{[j,n]:C(i^{\prime})<t\text{ for some }i^{\prime}\in[j,n],j>i\}$. For a prefix $Q^{-}\in Q_{C}^{-}(t)$ where $Q^{-}=[1,j]$, we let $\overline{Q^{-}}=[j+1,i-1]$, and similarly for a suffix $Q^{+}=[j,n]$, we let $\overline{Q^{+}}=[i+1,j-1]$.
@@
Fix a token $t$ with $C^{-1}(t)=v$. First, suppose that there are tokens on both sides of $t$ that wish to cross it. That is, there are vertices $i<v$ with $C(i)>t$ and $j>v$ with $C(j)<t$. We note that $\Phi_{C}(t)$ counts tokens, and the only tokens we may count lie between (inclusively) the smallest choice of $i$ and the largest choice of $j$. Since $t$ also indicates the target vertex of token $t$, the smallest $i$ can be is $t-d_{max}+1$; any smaller, and there would be a token more than a distance of $d_{max}$ from its destination. Similarly, the largest $j$ can be is $t+d_{max}-1$. There are a total of $12d_{max}-$ 1 tokens in this range, and we may count all but one (the token $t$ itself) as part of $\Phi_{C}(t)$. So $\Phi_{C}(t)\leq 2d_{max}-2$.
@@
If no tokens wish to cross $t$, then $\Phi_{C}(t)=0$, and the result is trivially true. In all cases, we have that $\Phi_{C}(t)\leq 2d_{max}-1$. ∎